May 1997 (session 2)
To give feedback on this experiment, click
Westerbork (tony foley at Tue Jun 3 17:20:30 METDST 1997) Minor failures
Lost passes 3 and 4 because of
a spurious vacuum loss error message
on the field system.
In this experiment only RT5 (the one actually in
use for 18cm VLBI) is controlled
by the field system.
Onsala (A. Polatidis at Wed Jun 4 10:24:00 METDST 1997) Success
No known problems
Torun (K.Borkowski at Wed Jun 4 13:55:55 METDST 1997) Minor failures
First 2 minutes with wrong LO setting.
Effelsberg (dgraham at Thu Jun 5 08:17:52 METDST 1997) Success
Jodrell (Alastair Gunn at Thu Jun 5 10:53:17 METDST 1997) Success
Extreme RFI during this experiment. No other known problems.
Cambridge (Alastair Gunn at Thu Jun 5 12:45:59 METDST 1997) Success
No known problems.
Westerbork (tony foley at Mon Jun 9 14:27:55 METDST 1997) Minor failures
Also lost passes 11 and 19, because of same
false alarm about tape not having vacuum
Noto (C.Trigilio at Mon Jun 9 16:03:52 METDST 1997) Success
Medicina (Venturi at Tue Jun 10 16:02:04 METDST 1997) Minor failures
Until 12:44 UT local oscillator value was wrong. The experiment after
Urumqi (Li Suqin at Thu Jun 12 07:19:24 METDST 1997) Success
It is claudy , the colck offset =4.5+-0.5us.
No problem found.
Simeiz (A.Volvatch/LIG at Wed Jun 18 09:03:02 METDST 1997) Success
No problem reported, tapes shipped on 06.06.97.
Medicina (Fredrik T Rantakyro at Mon Jul 21 11:18:44 METDST 1997) Minor failures
The ANTAB file for the experiment can be found on VLBEER in the MAY97
directory with the name GL20.ANTAB. The file is ready to be used
with the AIPS tasks ANTAB/APCAL. Read the ANTAB file for any specific
comments regarding this experiment.
MPIfR (Scott Aaron at Mon Jun 29 11:38:26 METDST 1998) Minor failures
Correlation completed. Hancock observed with the wrong frequencies
during the entire session, so no data were correlated. The tapes
from Jodrell were of low quality, especially in reverse passes.